OpenStudy (anonymous):

f(x)=mx+c why does an inverse function not exist if m=0

OpenStudy (anonymous):

i understand how g(x)=y/0 -c/0

OpenStudy (anonymous):

Well let m=0 f(x)=0*x+c =0+c=c So then f(x)=c

OpenStudy (anonymous):

Fine lets do it like that y=0*x+C Now lets fine the inverse: \[x=0*y+c\] \[y=\frac{x-c}{0}\]

OpenStudy (er.mohd.amir):

inverse is not define because it give infinity when m=0.

OpenStudy (anonymous):

@Maretch Do you understand?